can any one help me find the slope for each line? 1. y= -3/4+7 m= 2. x=y m= 3. 3x-8y=4 m=
Answer by Alan3354(31505)

1. y= -3/4+7 m=
2. x=y m=
3. 3x-8y=4 m=
----------
1. Did you mean y= -(3/4)x+7 ? There's no x in yours.
2. put in slope-intercept form, y=mx+b
y=x+0
m = 1
3. 3x-8y=4
-8y = 4-3x
8y=3x-4
y = (3/8)x -1/2
m = 3/8